A reservation clerk worked 12.6 hrs one day. She spent twice as much time entering new reservations as she did verifying old ones and one and a half as much time calling to confirm reservations as verifying old ones. How much time did she spend entering new reservations?
Thank you in advance
Chris Answer by htmentor(1343) 
 A reservation clerk worked 12.6 hrs one day. She spent twice as much time entering new reservations as she did verifying old ones and one and a half as much time calling to confirm reservations as verifying old ones. How much time did she spend entering new reservations?
======================================================
Let n = the time spent on new reservations
Then the time spent on old reservations is n/2
The time spent confirming reservations is 3/2 the time spent on old reservations
But the time spent on old reservations is n/2
Therefore, we can express the time spent confirming reservations as (3/2)(n/2)=3n/4
So the total time spent, expressed in terms of n, is:
n + n/2 + 3n/4 = 12.6
Solve for n:
(4n + 2n + 3n)/4 = 12.6
9n = 50.4
n = 5.6
So the time spent on new reservations is 5.6 hrs